# Inter-Satellite Link Configuration for Fast Delivery in Low-Earth-Orbit Constellations

### v1

@@ 1 Introduction
Recent low-Earth-orbit (LEO) constellations such as Starlink [fcc21], OneWeb [fcc20oneweb], Amazon’s Kuiper [fcc20kuiper], and AST SpaceMobile [fcc25ast_mod] aim to provide global connectivity by providing worldwide coverage. Well-configured optical inter-satellite links (ISLs) can enhance the efficiency of these constellations, making them appealing for latency-critical services like real-time remote control, remote surgery and telecare, high-frequency trading, blockchain synchronization, and emergency alerting [bozkurt2017why]. In free space, light travels approximately 300 km per millisecond (covering the geodesic distance between the North and South Poles in about 66.7 ms), roughly 47% faster than in fiber [handley2018delay]. Hence, a well-designed LEO ISL network can potentially surpass the fastest terrestrial cables on intercontinental routes. Achieving such capabilities requires that the worst-case end-to-end latency be tightly bounded.
@@
Meeting these latency requirements relies on the unique geometry of LEO constellations. Geostationary Earth Orbit (GEO) satellites, positioned at approximately $35{,}786\;\text{km}$, incur round-trip propagation delays of hundreds of milliseconds, far too high for interactive or low-latency applications. In contrast, LEO satellites, orbiting below $2{,}000\;\text{km}$, enable tightly interconnected optical meshes with per-hop propagation delays an order of magnitude smaller than those in GEO constellations. At the same time, LEO networks’ launch cost and transmit power requirements are more favorable than those of medium-Earth-orbit (MEO) systems. This positions LEO constellations as the premier platform for low-latency global networking in the foreseeable future [ntrs_2010_mooring_nasa, berry2024spectrum].
@@
Unlike terrestrial networks that use distributed protocols such as Internet Protocol (IP) and Border Gateway Protocol (BGP), LEO constellations operated by a single service provider, such as Starlink, can employ centralized routing strategies. Routes and forwarding tables can be pre-computed for short snapshots during which satellite positions and candidate links are treated as static, then pushed to all nodes before the next snapshot begins [handley2018delay]. This determinism eliminates the need for in-orbit route discovery and enables cut-through switching [kermani1979virtual], or multiprotocol label switching (MPLS)-style label forwarding [rfc3031], allowing each packet to be steered with microsecond-scale processing overhead per hop [heine2023end]. Moreover, if ISL capacity is provisioned to handle the offered traffic demands, queuing delays at each node can also be constrained to the microsecond scale with high probability. Under these conditions, where per-hop processing delay is negligible compared to propagation delay, end-to-end latency is dominated by the cumulative propagation delay across all traversed links. Since each hop introduces an additional propagation segment, the total delay grows approximately in proportion to the number of hops taken along the path. While minimizing the absolute path distance would be ideal, we instead use the network diameter—the largest shortest-path cost between any two satellites—as a structural proxy for latency. While a general problem is formulated using arbitrary (positive) link cost, we use hop count as the specific measure for path cost to solve the optimization problem in simulations. This choice is justified because LEO constellations form sparse geographical graphs (each satellite has only a few inter-satellite links), and link lengths are upper-bounded by orbital geometry. Consequently, a lower hop count typically implies a straighter, more direct route, reducing cumulative propagation distance and avoiding inefficient zig-zagging across the network.
@@
Reducing the diameter is straightforward in principle: add more inter-plane ISLs and choose them so that distant orbits are bridged early in a path. In practice, however, each satellite is limited to a small number of steerable laser terminals, and any candidate link must satisfy distance and line-of-sight constraints not just at one instant but ideally throughout an orbital period. The design problem is therefore to select a sparse, physically feasible set of inter-plane ISLs that keeps the diameter small while remaining stable over time. This paper addresses this challenge for a Walker–Delta constellation where each satellite maintains two fixed intra-plane links and may activate up to two inter-plane links (as in Starlink) [fcc21, wang1993structural]. The ISL-selection task is cast as a constrained diameter-minimization problem, and an iterative local-search algorithm is proposed to refine an initial feasible topology. Extensive simulations show that the resulting configurations (i) respect all geometric and hardware limits, (ii) remain viable across multiple orbital periods, and (iii) reduce worst-case hop count and thus latency. By quantifying the trade-off between hop reduction and long-term link stability, the study offers concrete guidance for next-generation LEO network deployments.
@@ 2 Related Work
The use of large-scale LEO satellite constellations for global broadband coverage is proposed as a promising direction for enabling low-latency, scalable communication. In such systems, satellites are deployed in dense orbital configurations and interconnected through ISLs to form a mesh-like network [handley2018delay]. It is argued that the effectiveness of these constellations depends heavily on how the underlying inter-satellite topology is designed, with efficient link assignment and routing considered essential for supporting time-sensitive applications.
@@
Several strategies for ISL topology design have been explored in prior work. A time-division-based method for ISL topology generation has been proposed, in which different sets of feasible links are periodically activated to maintain long-term connectivity [chu2018time]. Focusing on the ISL assignment problem for navigation constellations, link scheduling has been optimized for time synchronization and autonomous orbit determination. To solve this multi-objective problem under payload and visibility constraints, assignment algorithms based on satellite layering and weighting within a discrete-time system of superframes have been developed [sun2022inter]. To tackle the computational complexity of the large-scale ISL scheduling problem, it has been modeled as a time-discrete network multi-commodity flow problem, and a data-driven parallel adaptive large neighborhood search (DP–ALNS) algorithm has been proposed to find solutions that minimize total data transmission delay [liu2022data]. These approaches emphasize temporal reachability but do not explicitly minimize worst-case broadcast latency or network diameter.
@@
Multicast and broadcast strategies in LEO constellations have also been explored, including core-based multicast trees and disruption-tolerant mechanisms for multilayered satellite networks [cheng2007core, zheng2010routing]. The broader challenge of routing in LEO networks has been extensively surveyed, with trade-offs identified between centralized and distributed protocols and an emphasis placed on the need for algorithms that can adapt to predictable topological changes [xiaogang2016survey]. Building on the predictable topology of LEO constellations, a software-defined networking (SDN) based approach has been introduced in which a centralized terrestrial controller pre-computes routing tables for all anticipated topologies. These routing tables are then distributed to space nodes in advance, enabling them to switch based on time or network context, thereby eliminating in-network routing convergence time and reducing the computational load on satellites [corici2024sdn]. While these methods improve resilience and group communication, they often rely on idealized routing assumptions or neglect geometric constraints inherent to LEO dynamics.
@@
Other efforts have focused on low-latency path design under stochastic satellite distributions. Tools from stochastic geometry have been used to model spatial connectivity and evaluate average-case latency [wang2022stochastic], and congestion-aware routing portocols have been developed to improve throughput under dynamic traffic conditions [dai2021distributed]. However, these approaches generally target statistical performance rather than deterministic guarantees or worst-case latency bounds.
@@
Further analysis of path selection has introduced a theoretical model for explicitly estimating the ISL hop-count between ground users, showing how the hop-count is determined by user latitudes and longitude differences [chen2021analysis]. An explicit analytic algorithm was later proposed to solve the shortest distance path (SDP) problem by modeling it with satellite phase and converting it into a total phase offset problem that can be solved directly without graph-based iterations. It has also been proven that almost all SDPs lie within the minimum hop path (MHP) set, thereby simplifying the search space for optimal routing [chen2024shortest].
@@ 3 System Model
We model the network using a static snapshot of satellite positions. This approach allows routes to be precomputed for each epoch, enabling deterministic path planning and reducing in-network operational overhead. Such a pre-planned environment is ideally suited for low-latency forwarding techniques like cut-through switching and MPLS. In cut-through switching, a packet is forwarded to the next satellite as soon as the destination address is read, without the full packet being received [kermani1979virtual]. This minimization of buffering delay is particularly effective in high-speed optical networks where propagation delay dominates. Similarly, MPLS enables efficient packet forwarding by attaching short labels to packets, allowing each satellite to determine the next hop based on precomputed paths rather than full network lookups [rfc3031]. These features align naturally with the centralized control and predictable topology of LEO satellite networks.
Building on this snapshot-based framework, we consider two models for inter-plane link feasibility that reflect different operational assumptions. In the first, a link is considered feasible if it satisfies geometric constraints, namely distance and line-of-sight, at a single snapshot in time. This snapshot-only model offers greater flexibility in link selection but does not guarantee long-term physical validity as satellites move along their orbits. To address, this we consider a viability-constrained model that requires a link to remain feasible throughout multiple orbital cycles. This stricter condition yields topologies that are more robust and operationally deployable, as they reduce the need for frequent link reconfiguration. This distinction is particularly important in practice, as reconfiguring optical ISLs is non-trivial. Switching from one target to another requires physically steering laser terminals, re-acquiring beam alignment, and establishing a stable optical connection. Recent studies have shown that the setup time for such links can be on the order of several seconds, due to the delays associated with beam acquisition and pointing adjustments [bhattacharjee2023laser]. These delays introduce non-negligible communication gaps and limit the feasibility of frequent link switching. As a result, topologies that remain valid over extended time periods are preferred for maintaining stable and low-latency performance in LEO satellite networks.
@@ 6 Simulation & Results
The simulation is conducted using the parameters specified in Table 1. These parameters are inspired by Starlink Shell 1 configurations [fcc21]. The optimization runs for $n=300$ iterations per trial. A degree reinforcement phase is triggered every $K=15$ iterations, while in all other steps, the random replacement phase modifies a subset of $m=20$ satellites. This entire process is repeated across 50 randomized initializations to ensure robustness.
@@
The performance of each graph is evaluated through an all-source shortest-path analysis to determine its key latency metrics. A breadth-first search (BFS) is executed from each satellite as a source node to find its longest shortest-path distance in hops [kozen1992depth]. The network diameter and average maximum hop count are then derived from this complete set of measurements.

### v2

@@ 1 Introduction
Recent low-Earth-orbit (LEO) constellations such as Starlink [12], OneWeb [11], Amazon’s Kuiper [10], and AST SpaceMobile [13] aim to provide global connectivity by providing worldwide coverage. Well-configured optical inter-satellite links (ISLs) can enhance the efficiency of these constellations, making them appealing for latency-critical services like real-time remote control, remote surgery and telecare, high-frequency trading, blockchain synchronization, and emergency alerting [3, 20]. In free space, light travels approximately 300 km per millisecond (covering the geodesic distance between the North and South Poles in about 66.7 ms), roughly 47% faster than in fiber [14]. Hence, a well-designed LEO ISL network can potentially surpass the fastest terrestrial cables on intercontinental routes. Achieving such capabilities requires that the worst-case end-to-end latency be tightly bounded.
@@
Meeting these latency requirements relies on the unique geometry of LEO constellations. Geostationary Earth Orbit (GEO) satellites, positioned at approximately $35{,}786\;\text{km}$, incur round-trip propagation delays of hundreds of milliseconds, far too high for interactive or low-latency applications. In contrast, LEO satellites, orbiting below $2{,}000\;\text{km}$, enable tightly interconnected optical meshes with per-hop propagation delays an order of magnitude smaller than those in GEO constellations. At the same time, LEO networks’ launch cost and transmit power requirements are more favorable than those of medium-Earth-orbit (MEO) systems. This positions LEO constellations as the premier platform for low-latency global networking in the foreseeable future [16, 1].
@@
Unlike terrestrial networks that use distributed protocols such as Internet Protocol (IP) and Border Gateway Protocol (BGP), LEO constellations operated by a single service provider, such as Starlink, can employ centralized routing strategies. Routes and forwarding tables can be pre-computed for short snapshots during which satellite positions and candidate links are treated as static, then pushed to all nodes before the next snapshot begins [14]. This determinism eliminates the need for in-orbit route discovery and enables cut-through switching [17], or multiprotocol label switching (MPLS)-style label forwarding [22], allowing each packet to be steered with microsecond-scale processing overhead per hop [15]. Moreover, if ISL capacity is provisioned to handle the offered traffic demands, queuing delays at each node can also be constrained to the microsecond scale with high probability. Under these conditions, where per-hop processing delay is negligible compared to propagation delay, end-to-end latency is dominated by the cumulative propagation delay across all traversed links. Since each hop introduces an additional propagation segment, the total delay grows approximately in proportion to the number of hops taken along the path. While minimizing the absolute path distance would be ideal, we instead use the network diameter—the largest shortest-path cost between any two satellites—as a structural proxy for latency. While a general problem is formulated using arbitrary (positive) link cost, we use hop count as the specific measure for path cost to solve the optimization problem in simulations. This choice is justified because LEO constellations form sparse geographical graphs (each satellite has only a few inter-satellite links), and link lengths are upper-bounded by orbital geometry. Consequently, a lower hop count typically implies a straighter, more direct route, reducing cumulative propagation distance and avoiding inefficient zig-zagging across the network.
@@
Reducing the diameter is straightforward in principle: add more inter-plane ISLs and choose them so that distant orbits are bridged early in a path. In practice, however, each satellite is limited to a small number of steerable laser terminals, and any candidate link must satisfy distance and line-of-sight constraints not just at one instant but ideally throughout an orbital period. The design problem is therefore to select a sparse, physically feasible set of inter-plane ISLs that keeps the diameter small while remaining stable over time. This paper addresses this challenge for a Walker–Delta constellation where each satellite maintains two fixed intra-plane links and may activate up to two inter-plane links (as in Starlink) [12, 24]. The ISL-selection task is cast as a constrained diameter-minimization problem, and an iterative local-search algorithm is proposed to refine an initial feasible topology. Extensive simulations show that the resulting configurations (i) respect all geometric and hardware limits, (ii) remain viable across multiple orbital periods, and (iii) reduce worst-case hop count and thus latency. By quantifying the trade-off between hop reduction and long-term link stability, the study offers concrete guidance for next-generation LEO network deployments.
@@ 2 Related Work
The use of large-scale LEO satellite constellations for global broadband coverage is proposed as a promising direction for enabling low-latency, scalable communication. In such systems, satellites are deployed in dense orbital configurations and interconnected through ISLs to form a mesh-like network [14]. It is argued that the effectiveness of these constellations depends heavily on how the underlying inter-satellite topology is designed, with efficient link assignment and routing considered essential for supporting time-sensitive applications.
@@
Several strategies for ISL topology design have been explored in prior work. A time-division-based method for ISL topology generation has been proposed, in which different sets of feasible links are periodically activated to maintain long-term connectivity [7]. Focusing on the ISL assignment problem for navigation constellations, link scheduling has been optimized for time synchronization and autonomous orbit determination. To solve this multi-objective problem under payload and visibility constraints, assignment algorithms based on satellite layering and weighting within a discrete-time system of superframes have been developed [23]. To tackle the computational complexity of the large-scale ISL scheduling problem, it has been modeled as a time-discrete network multi-commodity flow problem, and a data-driven parallel adaptive large neighborhood search (DP–ALNS) algorithm has been proposed to find solutions that minimize total data transmission delay [19]. These approaches emphasize temporal reachability but do not explicitly minimize worst-case broadcast latency or network diameter.
@@
Multicast and broadcast strategies in LEO constellations have also been explored, including core-based multicast trees and disruption-tolerant mechanisms for multilayered satellite networks [6, 26]. The broader challenge of routing in LEO networks has been extensively surveyed, with trade-offs identified between centralized and distributed protocols and an emphasis placed on the need for algorithms that can adapt to predictable topological changes [21]. Building on the predictable topology of LEO constellations, a software-defined networking (SDN) based approach has been introduced in which a centralized terrestrial controller pre-computes routing tables for all anticipated topologies. These routing tables are then distributed to space nodes in advance, enabling them to switch based on time or network context, thereby eliminating in-network routing convergence time and reducing the computational load on satellites [8]. While these methods improve resilience and group communication, they often rely on idealized routing assumptions or neglect geometric constraints inherent to LEO dynamics.
@@
Other efforts have focused on low-latency path design under stochastic satellite distributions. Tools from stochastic geometry have been used to model spatial connectivity and evaluate average-case latency [25], and congestion-aware routing portocols have been developed to improve throughput under dynamic traffic conditions [9]. However, these approaches generally target statistical performance rather than deterministic guarantees or worst-case latency bounds.
@@
Further analysis of path selection has introduced a theoretical model for explicitly estimating the ISL hop-count between ground users, showing how the hop-count is determined by user latitudes and longitude differences [4]. An explicit analytic algorithm was later proposed to solve the shortest distance path (SDP) problem by modeling it with satellite phase and converting it into a total phase offset problem that can be solved directly without graph-based iterations. It has also been proven that almost all SDPs lie within the minimum hop path (MHP) set, thereby simplifying the search space for optimal routing [5].
@@ 3 System Model
We model the network using a static snapshot of satellite positions. This approach allows routes to be precomputed for each epoch, enabling deterministic path planning and reducing in-network operational overhead. Such a pre-planned environment is ideally suited for low-latency forwarding techniques like cut-through switching and MPLS. In cut-through switching, a packet is forwarded to the next satellite as soon as the destination address is read, without the full packet being received [17]. This minimization of buffering delay is particularly effective in high-speed optical networks where propagation delay dominates. Similarly, MPLS enables efficient packet forwarding by attaching short labels to packets, allowing each satellite to determine the next hop based on precomputed paths rather than full network lookups [22]. These features align naturally with the centralized control and predictable topology of LEO satellite networks.
Building on this snapshot-based framework, we consider two models for inter-plane link feasibility that reflect different operational assumptions. In the first, a link is considered feasible if it satisfies geometric constraints, namely distance and line-of-sight, at a single snapshot in time. This snapshot-only model offers greater flexibility in link selection but does not guarantee long-term physical validity as satellites move along their orbits. To address, this we consider a viability-constrained model that requires a link to remain feasible throughout multiple orbital cycles. This stricter condition yields topologies that are more robust and operationally deployable, as they reduce the need for frequent link reconfiguration. This distinction is particularly important in practice, as reconfiguring optical ISLs is non-trivial. Switching from one target to another requires physically steering laser terminals, re-acquiring beam alignment, and establishing a stable optical connection. Recent studies have shown that the setup time for such links can be on the order of several seconds, due to the delays associated with beam acquisition and pointing adjustments [2]. These delays introduce non-negligible communication gaps and limit the feasibility of frequent link switching. As a result, topologies that remain valid over extended time periods are preferred for maintaining stable and low-latency performance in LEO satellite networks.
@@ 6 Simulation & Results
The simulation is conducted using the parameters specified in Table 1. These parameters are inspired by Starlink Shell 1 configurations [12]. The optimization runs for $n=300$ iterations per trial. A degree reinforcement phase is triggered every $K=15$ iterations, while in all other steps, the random replacement phase modifies a subset of $m=20$ satellites. This entire process is repeated across 50 randomized initializations to ensure robustness.
@@
The performance of each graph is evaluated through an all-source shortest-path analysis to determine its key latency metrics. A breadth-first search (BFS) is executed from each satellite as a source node to find its longest shortest-path distance in hops [18]. The network diameter and average maximum hop count are then derived from this complete set of measurements.